# José Luis Blanco-Claraco

### v1

@@ Abstract
Light Detection and Ranging (LiDAR)-based simultaneous localization and mapping (SLAM) is a core technology for autonomous vehicles and robots. Despite the intense research activity in this field, each proposed system uses a particular sensor post-processing pipeline and a single map representation format. The present work aims at introducing a revolutionary point of view for 3D LiDAR SLAM and localization: (1) using view-based maps as the fundamental representation of maps (“simple-maps”), which can then be used to generate arbitrary metric maps optimized for particular tasks, e.g. obstacle avoidance, real-time localization; and (2) by introducing a new framework in which mapping pipelines can be defined without coding, defining the connections of a network of reusable blocks much like deep-learning networks are designed by connecting layers of standardized elements. Moreover, the idea of including the current linear and angular velocity vectors as variables to be optimized within the Iterative Closest Point (ICP) loop is also introduced, leading to superior robustness against aggressive motion profiles without an IMU. The presented open-source ecosystem, released to ROS 2 and whose core component is dubbed MOLA-LO (MOLA LiDAR odometry), includes tools and prebuilt pipelines covering all the way from data acquisition to map editing and visualization, real-time localization, loop-closure detection, or map georeferencing from consumer-grade Global Navigation Satellite System (GNSS) receivers. Extensive experimental validation reveals that the proposal compares well to, or improves, former state-of-the-art (SOTA) LiDAR odometry systems, while also successfully mapping some hard sequences where others diverge. A proposed self-adaptive configuration has been used, without parameter changes, for all 3D LiDAR datasets with sensors between 16 and 128 rings, and has been extensively tested on 83 sequences over more than 250 km of automotive, hand-held, airborne, and quadruped LiDAR datasets, both indoors and outdoors, with single or multiple simultaneous LiDAR scanners. Additional specialized pipelines are also provided for 2D LiDARs and for drones, as examples of the framework flexibility. The possibility of handling a larger variety of robotics datasets also becomes a valuable tool in itself, easing benchmarking SLAM and perception solutions. The open-sourced implementation is available online at https://github.com/MOLAorg/mola.
@@ 1 Introduction
3D LiDAR is among the most widely used sensors in contemporary mobile robots, autonomous vehicles, and Unmanned Aerial Vehicles (UAV) (Elhousni and Huang (2020); Lee et al. (2023)). While vision and RADAR are also relevant technologies and may become essential for many tasks, LiDAR will likely remain a core sensor for years to come (Roriz et al. (2022)) due to a number of intrinsic advantages: direct and accurate 3D sensing with wide field of view, robustness against varying sunlight conditions, reduced computation demand to obtain dense point clouds, etc.
@@
Among all the potential applications of LiDAR sensing, this work focuses on the problems of pose tracking and mapping. We call localization to pose tracking on a prebuilt environment model that remains unmodified. Positioning in unknown environments becomes the well-known Simultaneous Localization and Mapping (SLAM) problem (Cadena et al. (2016)). When only the most recently-observed parts of the environment are kept, the term odometry is used instead, hence we will leave the category SLAM for those approaches that are aware of loop closures (Grisetti et al. (2010)). That is, odometry methods provide local accurate estimation of relative motion in the short term, while SLAM must also ensure global consistency of both, the trajectory and the world model.
@@
Therefore, when LiDAR is the unique sensor in an odometry system it is called LiDAR odometry (LO). The main component presented in this work falls into that category. LO can be used alone or as part of a larger system, leading to full metric SLAM (e.g. Cartographer in Hess et al. (2016)) or multi-robot SLAM (e.g. multi S-Graphs in Fernandez-Cortizas et al. (2024)) systems. In the literature, we can find how additional sensors have been successfully integrated with LO, leading to LiDAR Inertial odometry (LIO) (Shan et al. (2020)), Visual Inertial Odometry (VIO) (Leutenegger et al. (2015); Forster et al. (2016)), or Visual Lidar-Inertial Odometry (VLIO) (Shan et al. (2021)). However, we will show that, for most conditions found in practice, our system is able to achieve good performance and accuracy relying on LiDAR alone, notwithstanding that even better results might arise from multimodal sensor fusion.
@@
The present work defends that LO and SLAM systems can be built in a way that maximizes the reusability of their components, making them more flexible and easier to modify by end users, while also allowing resulting world models (“maps”) to be built minimizing the loss of information with respect to raw sensor data. Two core ideas arise from these goals: (i) odometry and SLAM pipelines, together with their associated data structures, should be refactored into their smallest reusable components, allowing building whole systems from scratch by wiring them together and setting their parameters much like Deep Neural Networks (DNN) are designed by stacking prebuilt elements; (ii) disregarding the specific map type or features used by a particular LO or SLAM system, the most versatile map data structure is a view-based map (dubbed simplemaps in our framework), where a sparse set of key-frames are retained together with their fundamental information, e.g. raw sensor observations, kinematic state of the vehicle, etc. The first usage of the term view-based SLAM can be found in the 2000s (Eustice et al. (2005); Konolige et al. (2010)), although the idea can be traced back to the ground-breaking work Lu and Milios (1997) which served as the seed for an extremely successful approach to mapping dubbed Graph-SLAM (Grisetti et al. (2010)). A more recent application of this idea of keeping the raw sensor information in graph-based SLAM can be found within Google’s Cartographer (Hess et al. (2016)) open-source implementation (refer to “assets writer”), although not mentioned in the original publication.
@@
Therefore, building upon the two ideas above, an open-source software ecosystem is provided to the community, whose components are enumerated in Appendix A. Note that none of the presented algorithms are based on learning methods, although the modular nature of the framework allows their integration as future works.
@@
The rest of the article is organized as follows. Section 2 reviews previous works closely related to the presented contributions. Next, the proposed architecture and their individual parts are introduced in detail in Section 3. Experimental validation is provided in three parts: Section 4 presents extensive quantitative performance metrics of the proposed SLAM system, section 5 illustrates in a qualitative way some use cases of our system, including situations and datasets where other SOTA systems diverge, and section 6 shows localization results with previously-built maps. We wrap up with conclusions in section 7.
@@ 2.1 LiDAR odometry (LO)
The goal of a LO system is taking successive LiDAR scans and estimating the vehicle pose changes in between them. At the core of most proposed systems there exist modified versions of the Iterative Closest Point (ICP) algorithm (Besl and McKay (1992)), which solves that nonlinear problem in an iterative fashion, finding successive approximations for the sough relative pose between two input point clouds, looking for potential point-to-point pairings at each iteration. That process is often called scan matching or cloud registration.
@@
We can extract four key findings from past related works. First of all, existing approaches can be classified according to whether they try to match the latest observation against other past individual observations (the scan-to-scan approach), or against an incrementally-built local map (the scan-to-model or scan-to-map approach). It has been experimentally found that scan-to-scan works extremely well for 2D LiDARs, whereas the uneven distribution of measured points in 3D LiDARs favors the scan-to-model approach, as demonstrated with recent successful works like Cartographer (Hess et al. (2016)), IMLS-SLAM (Deschaud (2018)), or KISS-ICP (Vizzo et al. (2023)). Following this line, the system proposed in this work hence also adheres to the scan-to-model design (see §3.5).
@@
Secondly, it is well known that trying to use ICP to register the raw points from a 3D LiDAR leads to extremely poor results. This can be explained by the uneven distribution of sampled points in the environment, with a higher density of near points that then dominate the cost function of the ICP optimization stage. This fact, together with the common use of LiDARs with a circular scanning pattern, leads to degenerate ICP solutions where nearby rings are aligned with nearby rings, disregarding the contribution from points sampled from distant parts of the environment. To overcome this problem, downsampling has been extensively used in the literature. One of the first successful LO systems, Zhang and Singh (2014), splits the input clouds into two distinct subsets (corners and planar patches), performing a subsampling to keep only a maximum of 2 edge or 4 planar points per planar scan. Recent works (Deschaud (2018); Vizzo et al. (2023)) employ voxel-based sampling, where raw points are first classified into 3D voxel data structures and then a single representative point is extracted from each occupied voxel following a given heuristic criterion (e.g. the average, the first point, etc.). LOCUS 2.0, introduced in Reinke et al. (2022), proposes an adaptive multi-level sampling of point clouds to try to maintain a constant number of points per scan. Following the same goal, another recent work dubbed SIMPLE (Bhandari et al. (2024)) also performs spatial subsampling, but implemented via KD-trees instead of voxels. Efficient sampling of point clouds for registration is a research field on its own. The libpointmatcher project, started with Pomerleau et al. (2013), implements different strategies, such as those proposed in Rusinkiewicz and Levoy (2001) or in Labussière et al. (2020). Latest works tend to use simpler sampling schemes, but we have experimentally verified that even tiny implementation differences in the way points are sampled (e.g. using different hash functions in unordered voxel containers for successive downsampling stages, which determines which is the first point to fall into a given voxel) have a relevant impact in the accuracy of the obtained trajectories. Therefore, our system offers simple voxel-based sampling schemes as the default choice, but it also offers other possibilities for the community to experiment and benchmark them, while also being easily extensible with new sampling algorithms (see §3.4), including those based on deep learning, e.g. following Lang et al. (2019).
@@
Thirdly, all modern approaches to LO have replaced the original closed-form least-squares formulation in the original ICP paper (Besl and McKay (1992)) with less efficient but more robust alternatives. In particular, instead of leaving the cost function to uniquely depend on the L2 distances between pointwise pairings, two enhancements are nowadays common: (i) employing robustified least squares, that is, robust kernel or loss functions (as in Deschaud (2018); Vizzo et al. (2023)), and (ii) using more geometric constraints apart of point-to-point pairings. On this latter point, the most common alternative is using point-to-plane pairings, as used in the original LOAM (Zhang and Singh (2014)) or, recently, in MAD-SLAM (Ferrari et al. (2024)). LOCUS 2.0 (Reinke et al. (2022)) introduces the estimated normals in the cost function, while another prominent work in this sense, MULLS (Pan et al. (2021)), includes several potential geometric pairings simultaneously. A word is in order about SIMPLE (Bhandari et al. (2024)), which employs a different approach: its cost function avoids the determination of point-to-point pairings by just adding Gaussian-like “rewards” for each nearby point, which can be seen as a one-to-many point pairing policy. Learning from all this body of past works, our framework provides a generic ICP-like optimization algorithm at its core, with user-extensible cost functions, predefined cost functions for point-to-point, point-to-plane or point-to-line (among others), different robust kernels, and the possibility of multiple correspondences (see §3.7).
@@
The fourth lesson learned from past works comes from KISS-ICP (Vizzo et al. (2023)), which demonstrated that dynamically adaptable parameters are beneficial for two reasons: (i) they can improve the overall accuracy since parameters adapt automatically as the environment or the motion profile change over time, and (ii) they reduce the need for manually tuning parameters. From that work we have adopted (and modified) their idea of using an adaptive threshold for determining pairings, and extended the idea to many other parameters along the entire processing pipeline (see §3.6).
@@ 2.2 LiDAR localization
Localization with a predefined, static map and a 3D LiDAR is a topic with less research activity than LO and SLAM, despite the fact that one of the fundamental reasons one may want to build a map is to use it for navigation, which implies the ability to keep the vehicle localized in the map frame of reference. As in any other estimation problem, we could split classic approaches (those not based on deep learning) into those relying on parametric or on non-parametric distributions. The latter typically involves using particle filtering (PF) (Thrun et al. (2001)), which are an excellent approach for vehicles constrained to trajectories on SE(2), but whose performance may degrade due to the curse of dimensionality for SE(3) motion. Still, it was shown in Blanco-Claraco et al. (2019) how it is possible to keep a vehicle well-localized in SE(3) using a PF, from raw 3D LiDAR data and wheels odometry, much faster than real time. This line of work was enhanced by detecting and tracking features (pole-like objects) in Schaefer et al. (2021), still using a PF. On the other hand, parametric methods for localization, such as Hess et al. (2016) or Lee and Ryu (2024), typically rely on an ICP-like optimization loop to find the best vehicle pose at a given instant.
Since we believe that both approaches have their applicability niches, both solutions are provided in the present framework (see §6).
@@ 2.3 Graph optimization
The problem of finding the global optimal for a set of rigid transformations between a sequence of key-frames is a very well studied one in the robotics community, either in its planar SE(2) or spatial SE(3) form. It has been dubbed Graph-SLAM (Grisetti et al. (2010)), pose-graph optimization (Tian et al. (2021)), or synchronization if mostly interested in the rotational parts (Wang and Singer (2013); Carlone et al. (2015)).
The goal of all these methods is to find the poses between the key-frames that minimize the mismatch of (typically sparse) relative pose observations. These techniques have an obvious direct application to optimization-based (Strasdat et al. (2012)) odometry frameworks that follow the frame-to-frame paradigm. Despite the present work follows the frame-to-map model instead, we mention graph optimization here since it becomes essential when handling loop closures.
The most relevant graph optimization libraries are reviewed next, with a focus on those tailored to graph SLAM rather than bundle adjustment. One of the earliest works in this field is probably TORO (Grisetti et al. (2007b, a, 2009)), which implemented efficient gradient descent for pose networks of arbitrary topology (multiple loop closures). Next, G2O was presented in Kümmerle et al. (2011), becoming extremely popular, to the point that it is still used nowadays in SOTA frameworks. Three G2O features that explain this success are: (i) its modular design that allowed for arbitrary graphs to be defined, including pose-only graphs and those required for visual bundle adjustment (Triggs et al. (2000)), (ii) the popularization of perturbation-based optimization on Lie groups (the $\boxplus$ notation) (Kümmerle et al. (2011); Sola et al. (2018); Blanco-Claraco (2021)), and (iii) the support for robust kernels (Chebrolu et al. (2021)), essential for outlier rejection.
Afterwards, two other popular frameworks were introduced: GTSAM (Dellaert (2021)), and CERES (Agarwal et al. (2023)). Both are frameworks for solving large, non-linear optimization problems, and both feature robust kernels and some degree of automated calculation of cost function derivatives. In our loop closure subsystem (described in §3.13) we chose GTSAM as the graph optimization back-end.
Figure 1: Main top-level data structures and procedural components of the presented framework: (a) odometry and (b) loop-closure subsystems, (c) metric map generation or update pipelines, and (d) a scheme of the contents of a metric map. Refer to discussion in §3.1. The contents of each subsystem are detailed in subsequent figures.

### v2

@@ Abstract
Light Detection and Ranging (LiDAR)-based simultaneous localization and mapping (SLAM) is a core technology for autonomous vehicles and robots. One key contribution of this work to 3D LiDAR SLAM and localization is a fierce defense of view-based maps (pose graphs with time-stamped sensor readings) as the fundamental representation of maps. As will be shown, they allow for the greatest flexibility, enabling the posterior generation of arbitrary metric maps optimized for particular tasks, e.g. obstacle avoidance, real-time localization. Moreover, this work introduces a new framework in which mapping pipelines can be defined without coding, defining the connections of a network of reusable blocks much like deep-learning networks are designed by connecting layers of standardized elements. We also introduce tightly-coupled estimation of linear and angular velocity vectors within the Iterative Closest Point (ICP)-like optimizer, leading to superior robustness against aggressive motion profiles without the need for an IMU. Extensive experimental validation reveals that the proposal compares well to, or improves, former state-of-the-art (SOTA) LiDAR odometry systems, while also successfully mapping some hard sequences where others diverge. A proposed self-adaptive configuration has been used, without parameter changes, for all 3D LiDAR datasets with sensors between 16 and 128 rings, and has been extensively tested on 83 sequences over more than 250 km of automotive, hand-held, airborne, and quadruped LiDAR datasets, both indoors and outdoors. The system flexibility is demonstrated with additional configurations for 2D LiDARs and for building 3D NDT-like maps. The framework is open-sourced online: https://github.com/MOLAorg/mola.
@@ 1 Introduction
3D LiDAR is among the most widely-used sensors in contemporary mobile robots, autonomous vehicles, and Unmanned Aerial Vehicles (UAV) (Elhousni and Huang,, 2020; Lee et al., 2023). While vision and RADAR are also relevant technologies and may become essential for many tasks, LiDAR will likely remain a core sensor for years to come (Roriz et al., 2022) due to a number of intrinsic advantages: direct and accurate 3D sensing with wide field of view, robustness against varying sunlight conditions, reduced computation demand to obtain dense point clouds, etc.
@@
Among all the potential applications of LiDAR sensing, this work focuses on the problems of pose tracking and mapping. We call localization to pose tracking on a prebuilt environment model that remains unmodified. Positioning in unknown environments becomes the well-known Simultaneous Localization and Mapping (SLAM) problem (Cadena et al., 2016). When only the most recently-observed parts of the environment are kept, the term odometry is used instead, hence we will leave the category SLAM for those approaches that are aware of loop closures (Grisetti et al., 2010). That is, odometry methods provide local accurate estimation of relative motion in the short term, while SLAM must also ensure global consistency of both, the trajectory and the world model.
@@
Therefore, when LiDAR is the unique sensor in an odometry system it is called LiDAR odometry (LO). The main component presented in this work falls into that category. LO can be used alone or as part of a larger system, leading to full metric SLAM, e.g. Cartographer in Hess et al. (2016), or multi-robot SLAM systems, e.g. multi S-Graphs in Fernandez-Cortizas et al. (2024). In the literature, we can find how additional sensors have been successfully integrated with LO, leading to LiDAR Inertial odometry (LIO) (Shan et al., 2020; Xu et al., 2022), Visual Inertial Odometry (VIO) (Leutenegger et al., 2015; Forster et al., 2016), or Visual Lidar-Inertial Odometry (VLIO) (Shan et al., 2021). However, we will show that, for most conditions found in practice, our system is able to achieve good performance and accuracy relying on LiDAR alone, notwithstanding that even better results might arise from multimodal sensor fusion.
@@
The present work defends that LO and SLAM systems can be built in a way that maximizes the reusability of their components, making them more flexible and easier to modify by end users, while also allowing resulting world models (“maps”) to be built minimizing the loss of information with respect to raw sensor data. Two core ideas arise from these goals: (i) odometry and SLAM pipelines, together with their associated data structures, should be refactored into their smallest reusable components, allowing building whole systems from scratch by wiring them together and setting their parameters much like Deep Neural Networks (DNN) are designed by stacking prebuilt elements; (ii) disregarding the specific map type or features used by a particular LO or SLAM system, the most versatile map data structure is a view-based map (dubbed “simple maps” in our framework), where a sparse set of key-frames are retained together with their fundamental information, e.g. raw sensor observations, kinematic state of the vehicle, etc. The first usage of the term view-based SLAM can be found in the 2000s (Eustice et al., 2005; Konolige et al., 2010), although the idea can be traced back to the ground-breaking work Lu and Milios (1997) which served as the seed for an extremely successful approach to mapping dubbed Graph-SLAM (Grisetti et al., 2010). A more recent application of this idea of keeping the raw sensor information in graph-based SLAM can be found within Google’s Cartographer (Hess et al., 2016) open-source implementation (refer to “assets writer”), although not mentioned in the original publication.
@@
Therefore, building upon the two ideas above, an open-source software ecosystem is provided to the community, whose components are enumerated in Appendix LABEL:apx:software. Note that none of the presented algorithms are based on learning methods, although the modular nature of the framework allows their integration as future works.
@@
Therefore, building upon the two ideas above, this work introduces an open-source software ecosystem to the community. Note that none of the presented algorithms are based on learning methods, although the modular nature of the framework allows their integration as future works. To sum up, the key contributions of this work are:
  • A method to easily integrate vehicle velocity as an unknown within ICP loops, allowing better point cloud undistortion, subject to a constant velocity assumption during each sensor sweep.
  • A simple method to let ICP take into account prior uncertainty, including that from a constant velocity motion model or other sources of odometry, e.g. wheel encoders or an IMU.
The rest of the article is organized as follows. Section 2 reviews previous works closely related to the presented contributions. Next, the proposed architecture and their individual parts are introduced in detail in Section 3. Experimental validation is provided in three parts: Section 4 presents extensive quantitative performance metrics of the proposed SLAM system, section 5 illustrates in a qualitative way some use cases of our system, including situations and datasets where other SOTA systems diverge, and section 6 shows localization results with previously-built maps. Furthermore, ablation studies are presented in section 7 and we wrap up with conclusions in section 8.
@@ 2.1 LiDAR odometry (LO)
The goal of a LO system is taking successive LiDAR scans and estimating the vehicle pose changes in between them. At the core of most proposed systems there exist modified versions of the Iterative Closest Point (ICP) algorithm (Besl and McKay, 1992), which solves that nonlinear problem in an iterative fashion, finding successive approximations for the sough relative pose between two input point clouds, looking for potential point-to-point pairings at each iteration. That process is often called scan matching or point cloud registration.
@@
We can extract four key findings from past related works. First of all, existing approaches can be classified according to whether they try to match the latest observation against other past individual observations (the scan-to-scan approach), or against an incrementally-built local map (the scan-to-model or scan-to-map approach). It has been experimentally found that scan-to-scan works extremely well for 2D LiDARs, whereas the uneven distribution of measured points in 3D LiDARs favors the scan-to-model approach, as demonstrated with recent successful works like Cartographer (Hess et al., 2016), IMLS-SLAM (Deschaud, 2018), or KISS-ICP (Vizzo et al., 2023). Following this line, the system proposed in this work hence also adheres to the scan-to-model design (see Section 3.4).
@@
Secondly, it is well known that trying to use ICP to register the raw points from a 3D LiDAR leads to extremely poor results. This can be explained by the uneven distribution of sampled points in the environment, with a higher density of near points that then dominate the cost function of the ICP optimization stage. This fact, together with the common use of LiDARs with a circular scanning pattern, leads to degenerate ICP solutions where nearby rings are aligned with nearby rings, disregarding the contribution from points sampled from distant parts of the environment. To overcome this problem, downsampling has been extensively used in the literature. One of the first successful LO systems, LOAM, introduced in Zhang and Singh (2014), splits the input clouds into two distinct subsets (corners and planar patches), performing a subsampling to keep only a maximum of 2 edge or 4 planar points per planar scan. This seminal work led to many variations, such as Lego-LOAM in Shan and Englot (2018) or PaGO-SLAM in Seo et al. (2022) which exploited an explicit ground plane segmentation. One of the most prominent derived works is Fast-LIO2 (Xu et al., 2022), which leverages tightly-coupled IMU-LiDAR integration with efficient data structures for fast and accurate mapping without the need to extract features. Other recent works (Deschaud, 2018; Vizzo et al., 2023) employ voxel-based sampling, where raw points are first classified into 3D voxel data structures and then a single representative point is extracted from each occupied voxel following a given heuristic criterion (e.g. the average, the first point, etc.). LOCUS 2.0, introduced in Reinke et al. (2022), proposes an adaptive multi-level sampling of point clouds to try to maintain a constant number of points per scan. Following the same goal, another recent work dubbed SIMPLE (Bhandari et al., 2024) also performs spatial subsampling, but implemented via KD-trees instead of voxels. Efficient sampling of point clouds for registration is a research field on its own. The libpointmatcher project, started with Pomerleau et al. (2013), implements different strategies, such as those proposed in Rusinkiewicz and Levoy (2001) or in Labussière et al. (2020). Latest works tend to use simpler sampling schemes, but we have experimentally verified that even tiny implementation differences in the way points are sampled (e.g. using different hash functions in unordered voxel containers for successive downsampling stages, which determines which is the first point to fall into a given voxel) have a relevant impact in the accuracy of the obtained trajectories. Therefore, our system offers simple voxel-based sampling schemes as the default choice, but it also offers other possibilities for the community to experiment and benchmark them, while also being easily extensible with new sampling algorithms (see Section 3.3), including those based on deep learning, e.g. following Lang et al. (2019).
@@
Thirdly, all modern approaches to LO have replaced the original closed-form least-squares formulation in the original ICP paper (Besl and McKay, 1992) with less efficient but more robust alternatives. In particular, instead of leaving the cost function to uniquely depend on the L2 distances between pointwise pairings, two enhancements are nowadays common: (i) employing robustified least squares, that is, robust kernel or loss functions as in (Deschaud, 2018; Chebrolu et al., 2021; Vizzo et al., 2023), and (ii) using more geometric constraints apart of point-to-point pairings. On this latter point, the most common alternative is using point-to-plane pairings, as used in the original LOAM (Zhang and Singh, 2014) or, more recently, in Fast-LIO2 (Xu et al., 2022) and in MAD-SLAM (Ferrari et al., 2024). LOCUS 2.0 (Reinke et al., 2022) introduces the estimated normals in the cost function, while another prominent work in this sense, MULLS (Pan et al., 2021), includes several potential geometric pairings simultaneously. A word is in order about SIMPLE (Bhandari et al., 2024), which employs a different approach: its cost function avoids the determination of point-to-point pairings by just adding Gaussian-like “rewards” for each nearby point, which can be seen as a one-to-many point pairing policy. Learning from all this body of past works, our framework provides a generic ICP-like optimization algorithm at its core, with user-extensible cost functions, predefined cost functions for point-to-point, point-to-plane or point-to-line (among others), different robust kernels, and the possibility of multiple correspondences (see Section 3.6).
@@
The fourth lesson learned from past works comes from KISS-ICP (Vizzo et al., 2023), which demonstrated that dynamically adaptable parameters are beneficial for two reasons: (i) they can improve the overall accuracy since parameters adapt automatically as the environment or the motion profile change over time, and (ii) they reduce the need for manually tuning parameters. From that work we have adopted (and modified) their idea of using an adaptive threshold for determining pairings, and extended the idea to many other parameters along the entire processing pipeline (see Section 3.5).
@@ 2.2 LiDAR localization
Localization with a predefined, static map and a 3D LiDAR is a topic with less research activity than LO and SLAM, despite the fact that one of the fundamental reasons one may want to build a map is to use it for navigation, which implies the ability to keep the vehicle localized in the map frame of reference. As in any other estimation problem, we could split classic approaches (those not based on deep learning) into those relying on parametric or on non-parametric distributions. The latter typically involves using particle filtering (Thrun et al., 2001), which are an excellent approach for vehicles constrained to trajectories on SE(2), but whose performance may degrade due to the curse of dimensionality for SE(3) motion. Still, it was shown in Blanco-Claraco et al. (2019) how it is possible to keep a vehicle well-localized in SE(3) using a particle filter, from raw 3D LiDAR data and wheels odometry, much faster than real time. This line of work was enhanced by detecting and tracking features (pole-like objects) in Schaefer et al. (2021), still using particle filtering. On the other hand, parametric methods for localization, such as Hess et al. (2016) or Lee and Ryu (2024), typically rely on an ICP-like optimization loop to find the best vehicle pose at a given instant.
Since we believe that both approaches have their applicability niches, both solutions are provided in the present framework (see Section 6).
@@ 2.3 Graph optimization
The problem of finding the global optimal for a set of rigid transformations between a sequence of key-frames is a very well studied one in the robotics community, either in its planar SE(2) or spatial SE(3) form. It has been dubbed Graph-SLAM (Grisetti et al., 2010, 2012), pose-graph optimization (Tian et al., 2021), or synchronization if mostly interested in the rotational parts (Wang and Singer, 2013; Carlone et al., 2015).
The goal of all these methods is to find the poses between the key-frames that minimize the mismatch of (typically sparse) relative pose observations. These techniques have an obvious direct application to optimization-based (Strasdat et al., 2012) odometry frameworks that follow the frame-to-frame paradigm. Despite the present work follows the frame-to-map model instead, we mention graph optimization here since it becomes essential when handling loop closures.
The most relevant graph optimization libraries are reviewed next, with a focus on those tailored to graph SLAM rather than bundle adjustment. One of the earliest works in this field is probably TORO (Grisetti et al., 2007b, a, 2009), which implemented efficient gradient descent for pose networks of arbitrary topology (multiple loop closures). Next, G2O was presented in Kümmerle et al. (2011), becoming extremely popular, to the point that it is still used nowadays in SOTA frameworks. Three G2O features that explain this success are: (i) its modular design that allowed for arbitrary graphs to be defined, including pose-only graphs and those required for visual bundle adjustment (Triggs et al., 2000), (ii) the popularization of perturbation-based optimization on Lie groups (the $\boxplus$ notation) (Kümmerle et al., 2011; Sola et al., 2018; Blanco-Claraco, 2021), and (iii) the support for robust kernels (Chebrolu et al., 2021), essential for outlier rejection.
Afterwards, two other popular frameworks were introduced: GTSAM (Dellaert, 2021), and CERES (Agarwal et al., 2022). Both are frameworks for solving large, non-linear optimization problems, and both feature robust kernels and some degree of automated calculation of cost function derivatives. In our loop closure subsystem (described in Section 3.12) we chose GTSAM as the graph optimization back-end.
Figure 1: Main top-level data structures and procedural components of the presented framework: (a) odometry and (b) loop-closure subsystems, (c) metric map generation or update pipelines, and (d) a scheme of the contents of a metric map. Refer to discussion in Section 3.1. The contents of each subsystem are detailed in subsequent figures.